Find the magnitude of the vector v=2i−5j+4k


1
Expert's answer
2021-10-26T10:20:46-0400

Magnitude of any vector is "|xi+yj+zk|=\sqrt{x^2+y^2+z^2}"

Hence the magnitude of this vector"=\sqrt{2^2+(-5)^2+4^2}"

"=\sqrt{45}"